Science Report – April 11th

MDRS Crew 192 Science Report (We are going to science the sh*t out of this!)

April 11, 2018 (SOL 3)

Richard Blakeman, Executive Officer

This crew has performed magnificently despite the challenges of having all the original science and engineering projects removed from the mission. The crew has pooled their individual and collective talents and shown incredible resourcefulness, creativity, imagination, and teamwork to develop multiple real-world science and engineering research and experiments.

The current list of projects that are being conducted include:

Spacesuit visor fogging study: Using off the shelf cleaning products which are used to develop mitigation strategies to solve the visor fogging issues. These are guaranteed to not scratch, damage or otherwise alter the visibility of the suits.

Hand exercises using hand relief, well-being balls: To measure comfort and hand dexterity during EVA

Crew wellness observations: Questionnaire designed to measure crew well-being during the duration of the mission

Crew weight measurements and analysis (EVA*): Pre and post weight for all personnel designed to measure body fluid loss during EVA

Crew muscle measurements: Physical measurements using tape measure to indicator of skeletal muscle glycogen reserve during mission

Ultrasonic rodent repulsion experiment: Using off the shelf, plug-in ultrasonic transmitter to observe if they deter rodent intrusions.

We noted a squeak at ~0830 today and have no confirmation but suspect there is a mouse about the house. We have also cleaned again and moved stove and frig to more thoroughly clean. What a mess behind both. Manual trap has been set IVO the squeak.

Astronomy discussions and visual observations since the main telescope is non-functional

Geology observations conducted during EVAs

EVA touch screen glove testing: We purchased several different kinds of touch pad sensitive gloves and anecdotally determining which ones are better for use with phone screen cameras.

Water contamination prevention and mitigation procedures

We have rolled the water tank from the trailer into the RAM and are going to do recurring EVAs to clean and sanitize the water tank for use with future crews to ease the requirement to make multiple trips into town for water saving fuel, vehicle wear and tear as well as cost of fuel for the society.

In honor of the achievements of Cosmonaut Yuri Gagarin, we have used a ration of our potatoes along with banana chips, apple slices and water. We heated this and combined it with yeast from the cupboard. The crew anticipates the ability to brew and enjoy a beverage made from distilled dehydrated potatoes on Yuri’s night (12 April).

The crew continues to take both still and video imagery for later analysis.

Sol Summary – April 11th

Crew 192 Sol Summary Report

Sol 3

Summary Title: Gone With the Wind

Author: Victoria Varone

Mission Status: Today was the first really hot day of our mission, a surprise in between the last few days of cool weather and wind. We continued to explore further out into the Martian terrain, exercising extra caution this time to remain hydrated and not push ourselves past our limits in hazardous conditions. The search for our missing science experiments goes on and we simultaneously improvise new science experiments to execute in their place.

Sol Activity Summary: Our day started early with our first EVA, where crew members Joe Dituri and Andreaa Radulescu went to explore another site to do a bit of hiking and geological investigation. When they returned we all had lunch together, which was followed by a brainstorming session for our next few EVAs.

Then we relaxed and watched The Martian, which was surprisingly similar to our own mission (missing crew member, same number of remaining crew members, affinity for duck tape), especially since not long after, a crazy wind storm kicked up.

Upon hearing a lot of commotion outside, we decided to have crew members Ashok Narayanamoorthi and Richard Blakeman head out on an engineering EVA prior to our scheduled evening EVA to check all around the hab. They found that a railing near the main door was broken because the front air lock door swung open and smashed railing. They made a couple of adjustments to secure the remaining piece as well as the hab door. Following that, the wind became so strong and started to present a safety hazard to the crew so the leader of the away team (Richard Blakeman) terminated the EVA.

Look Ahead Plan: We will continue to extend our range of Martian exploration as far as we possibly can, but some forecast hazardous weather has prompted us to change up some of our EVA plans. With strong winds predicted for tomorrow and Friday, we may choose to instead remain closer to the hab and work on our water tank cleanup project in the RAM and resume exploration over the weekend. Safety ALWAYS comes first.

Anomalies in Work: None

Weather: Hot and windy. The temperature reached approximately 83 degrees today and the wind storm was severe.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, GreenHab Report, EVA Request, EVA Report Journalist Report, Commander’s Report, Photo Report, Science Report

Support Requested: See Commander’s Report.

Commander Report – April 11th

Mission Commander report – Crew 192 Report SOL 3

This morning I made a modification to the scheduled EVA for health and comfort of the crew. I started an EVA ~30 minutes early intentionally to avoid the heat of the day which was (late information) projected to be hotter than expected. Please do not be upset with the crew. I assumed that I had the skill and ability to make minor modifications from leading multiple missions to myriad overseas destinations with significant consequence of failure. I did so fully understanding the EVAs and other non-critical mission parameters given my role and understanding all the reasons for the established rules such as crews possibly encountering those not in SIM. I know this now to be wrong and will adjust the EVA times to account for increased temperatures by scheduling earlier in the AM and later in the PM. I had no intention of making an already tenuous situation with the society worse by my actions… I apologize for my error and will ensure it does not happen again. Please accept my humble apologies. The crew is beginning to feel a “combative / oppressive dynamic” developing between external sources on our overall team and the society into which we may be being thrust. It is not our intention to be combative or disruptive …we intend only to maximize our learning and scientific potential while here. We have an will continue to build bridges while here and work towards peace.

Please understand that EVERY piece of correspondence that leaves this sight is reviewed and approved by me and altered with my …as well as every other crew member’s input as would be the case on Mars and in any other mission. There is no one man show on this team…please consider that when questioning why my name is not on a specific report.

On a lighter note we had Martian matinee today …using the time hot midday time to catch up on a movie because the evenings have been filled with planning.

LEARNED: The heat of the day is beginning to oppress our EVAs (cut morning EVA short due to temperature) so we will schedule EVAs earlier and later to coincide with cooler points of the day. Additionally, the weather is less predictable than initially thought.

IMPRESSED: I was impressed with the skill of Blue 04 (Victoria) who has been the go between and on all email correspondence and has been dealing with an internet speed just faster than glacial as well as minimal data capacity and still manages to upload myriad required reports. BZ Blue 04!

POA&M: Tomorrow the weather is supposed to turn to exceptionally high winds. High winds have started tonight. We have done a “batten the hatches down” EVA to look around the HAB in anticipation of the pending increased wind. Many items were removed from the HAB as requested by staff on site and should be removed before winds increase. We are planning a training and cleaning day tomorrow with EVAs very close to home.

~Dituri Sends (Blue 00)

Crew photos – April 10th

Sorry so late, this one was giving me trouble today.

Photo of the Day: HSO Ashok Narayanamoorthi on his first EVA on Mars

Sol Summary – April 10th

Crew 192 Sol 2 Summary Report

Sol 2

Summary Title: A Beautiful Day for EVAs

Author: Victoria Varone

Mission Status: It was a beautiful day for EVAs over here for Crew 192 on Mars. We continue to clean and organize things around the hab and have started getting more comfortable with EVAs, which will allow us to continue the search for fascinating natural wonders as well as our missing science experiments and crew member. The mission is going well.

Sol Activity Summary: The day started with a morning EVA, where three of our crew checked out the Marble Ritual site and surrounding area. They discovered some interesting geological structures and were particularly fascinated by a view of snow-capped mountains from a dusty, sandy red desert. After exploring, the crew returned to the hab, examined one of the hab’s water tanks for future cleaning, and headed inside for lunch.

After a great afternoon session of debriefing, brainstorming, and improvising for additional science experiments, the remaining crew members headed out for another EVA, coming back with excited descriptions of their hike in Kissing Camels Ridge. They also had some insight into the length of time of EVAs, noting that the amount of hiking they did in their 2.5 hour EVA expended just enough of their energy without causing serious dehydration. They’ve advise the rest of us to limit high-impact EVAs to about two and a half hours for safety.

We’re now resting and planning for tomorrow’s activities and may end the night with a space-based movie.

Look Ahead Plan: The goal is continue stretching our EVAs a little further as we keep searching for our missing science experiments and crew member. Based on tonight’s hiking EVA, we also plan to pay special attention to crew health on EVAs, making sure everyone is healthy and comfortable and no one becomes dehydrated and ill. A little attention is also being put towards possible activities for our Yuri’s Night celebration.

Anomalies in Work: None

Weather: The weather was beautiful again today, the skies started out clear and the weather cool, which worked well for the first EVA. It became overcast from the afternoon on, which allowed our second EVA crew members to explore the sites with different skies overhead.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, GreenHab Report, EVA Request, Journalist Report, Commander’s Report, Photo Report

Support Requested: None

EVA Report – April 10th

EVA Report

EVA #2 – 10Apr2018 – Sol 2

Crew members: Richard Blakeman, Ashok Narayanamoorthi, Victoria Varone

– Further explore Marble Ritual site
– Document EVA
– Examine third water tank for state of cleanliness and usability.

0815 Begin EVA
0830 Arrive at Marble Ritual
0900 Explore surrounding areas
1000 Return to hab site
1015 Examine third hab water tank
1030 Head back inside hab

Today was the crew’s second EVA of our mission. Three crew members (out of 5) went to explore a nearby site and surrounding areas, obtaining a lot of documentation of interesting geologic structures and formations and searching for possible signs of life in the reddish dirt. To head back to the hab, we traveled a little further with our vehicles to check out the landscape and utilized a turning point to turn back around. Upon arriving back at the hab, the rest of our crew inside requested that we take a quick survey of our third water tank to prepare it for being moved and cleaned. After that, we went back into the hab lunch.

1. Spirit Rover
2. Opportunity Rover
3. Blue ATV #3


EVA Report

EVA #3 – 10Apr2018 – Sol 2

Crew members: Joe Dituri, Andreaa Radulescu

– Explore a new site
– Search for missing science experiments

1630 – Entering the airlock
1635 – Move third water tank to RAM
1645 – Lock RAM and proceed to EVA
1730 – Arrive at Camel Ridge
1830 – Leave Camel Ridge
1900 – Arrive back at the hab

Today’s afternoon EVA was an opportunity to explore the Martian surface a little further than we did the day before. First, we completed a task for the crew, which was to move the third external water tank into our RAM for cleaning during our next few EVAs. After locking it inside the RAM, we headed out onto the Martian surface to an area called Kissing Camel Ridge, where we hiked for the majority of our EVA. We ended up finding a small badge of our missing crew member and decided to hike further in the area. The hike used up a lot of our energy, so we capped off our EVA at a little over two hours and headed back to the hab.

1. Red ATV
2. Blue ATV #3

Operations Report – April 10th

Crew 192 Operations Report 10Apr2018


Name of person filing report: Victoria Varone

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): 13 hours

Solar — SOC 100%

Diesel – 70%

Propane – 68% (telemetry reported) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – N/A

Water (trailer) – N/A
Water (static) – 1020 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133303.1

Toilet tank emptied: No. Last emptied April 9.

ATV’s Used: Blue #3, Red #1, and Blue #2.

Oil Added? No

ATV Fuel Used: 25%

# Hours the ATVs were Used today: 4

Notes on ATVs:

Deimos rover used: No
Hours: 0
Beginning charge: N/A
Ending charge: N/A
Currently charging: Yes

Spirit rover used: Yes
Hours: 2
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 2
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 0
Beginning charge: N/A
Ending charge: N/A
Currently charging: No

HabCar used and why, where? No

General notes and comments: A good day overall, the crew is doing very well.

Summary of internet: 569 MB remaining at 9:30PM.

Summary of suits and radios: Suits all charging, 4 one-piece suits used today. Radios working.

Summary of Hab operations: Changed out air filter, constructed minor and non-permanent communications modifications for radio clips in spacesuits, secured step outside of the side airlock.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: None.

Summary of RAM operations: The third large external water tank on the trailer was rolled into the RAM for cleaning during additional EVAs.

Summary of health and safety issues: None.

Questions, concerns, and requests to Mission Support: The hot water faucet in the hab kitchen appears to be faulty, it was difficult to shut off after use. To avoid the loss of water, we’re not going to use it.

Journalist Report – April 10th

Journalist Report – 10 April 2018

Author: Richard Blakeman

We started this morning on a high note as I led crew members Ashok Narayanamoorthi and Victoria Varone out on our first Martian EVA. We specifically scheduled the day’s first EVA in the morning hours to avoid the strong afternoon sun, as crew members Joe Dituri and Andreaa Radulesco reported getting a little hot during their first EVA yesterday afternoon.

This allowed us to stay out a bit longer and explore further into the area, where we documented many different geologic formations, colors, and structures. We were very excited to come across erosion that was likely caused by liquid water flow, as well as some interesting animals tracks, both of which suggest possible life on Mars.

Our second EVA took place late in the afternoon, where crew members Joe Dituri and Andreaa Radulesco began by moving our third large water tank into the RAM for cleaning, which will be continued tomorrow. They locked up the RAM and headed out to explore under an overcast, cloudy sky.

Today was also our first day conducting two EVAs where we alternated crew members to allow for the best possible use of energy and time and to allow the whole crew enough time to explore the Red Planet.

Crew morale remains outstanding as we’ve been finding many ways to occupy our time, including minor fixes and adjustments around the hab, experimenting with cooking our dehydrated supply of food, brainstorming EVAs, and playing games. We also just found the hab’s stash of movies which will supply a wonderful additional dose of entertainment for our movie-loving crew.

Blue Crew 192 will continue their adventures on Mars tomorrow.

Richard Blakeman

Greenhab Report – April 10th

GreenHab Report
Name: Andreea Radulescu
Date: 10 Apr 2018

Environmental control:

Ambient with door open
Shade cloth on
Fan on (15:15~18:30)

Average temperatures: 30.6℃
Low 14.6℃
High 45.7℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 5 gal

Time(s) of watering for crops: Once in the morning around 9am

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative: Picked few tomatoes and the crew ate them during brunch.

Support/supplies needed: None

Commander Report – April 10th

Mission Commander’s Report – SOL 2 – 10 April 2018

Author: Joe Dituri

With our morning EVA complete, we are formulating a plan for future EVAs and an effort to find our science suite and remnants of our science advisor. On our evening EVA, we were super excited to find a plaque card with our fallen science officer’s name on it. This renewed our optimism to find our science module in parts and salvage some of the science. The doctor continues to weigh the entire crew before and after EVA. So far most of the crew loses 1-3 lbs of body water per EVA depending upon duration. The doctor has continued to push fluids on the crew at all times for optimal health. He is also doing Skeletal muscle measurements for glycogen storage. We are performing isometric handgrip study prior to EVA to evaluate dexterity, numbness and comfort using EVA gloves. Yesterday we also started the well-being surveys. XO has started study to mitigate fogging of visors. Tomorrow the XO will submit an intermittent science report which will further detail science.

The crew played “Cards Against Humanity” last night. It served the purpose of lightening the mood after a hard day’s work. In the morning everyone woke refreshed and ready to start the day. The morning EVA served its purpose to indoctrination the three-person team for EVAs. The team explored the ridge to the east and found evidence of water as well as animal tracks.

The crew has worked up a plan to clean the water tank on the trailer. We took some of the EVA time to assess what was required to clean. We intend on rolling itinto the RAM tomorrow AM.

On the lighter side…some of the crew with longer hair compared dry shampoos and leave in conditioners and agreed to exchange product for a direct comparison taking into account a decreased ability to use water. Those participants who have insufficient hair to participate in the cross functional study will act as control group. Additionally hair constraint methods were discussed to reduce hair from being caught in the helmet

The crew took a great picture this morning with our red noses one and all to take a stand against childhood hunger. We were thinking about sending starving children some dehydrated food. Spirits are high and the crew is in a great mood.

LEARNED – We learned that the one-piece EVA suits have less mobility in the forward bending direction but they have different pressure points than the two-piece suit.

EVAs that involve a large amount of hiking should be limited to two or maybe two and a half hours unless there is supplemental water.

South of Kissing Camel Ridge, comms go out and you cannot get in touch with the hab.

IMPRESSED – I was impressed with the XO (Richard – Blue 01) as he led the EVA team and noticed the communication microphones were ill fitted. He devised a non – permanent alteration for the microphones which mimics a boom microphone and increased the effective communication. BRAVO ZULU to the XO.

POA&M – (Plan of Action and Milestones) – We are going to attempt to explore the Kissing Camel area where the science module plaque was found. We now have a training schedule for the crew to learn about decompression in space suits and psychological crew selection as well as radiation exposure and mitigation factors. Yuri’s night is the 12th of April and we made an attempt at planning some fun albeit limited.

~Dituri Sends (Blue 00)